# Data-Free Class-Incremental Gesture Recognition with Prototype-Guided Pseudo Feature Replay

### v1

@@ Abstract
Gesture recognition is an important research area in the field of computer vision. Most gesture recognition efforts focus on close-set scenarios, thereby limiting the capacity to effectively handle unseen or novel gestures. We aim to address class-incremental gesture recognition, which entails the ability to accommodate new and previously unseen gestures over time. Specifically, we introduce a Prototype-Guided Pseudo Feature Replay (PGPFR) framework for data-free class-incremental gesture recognition. This framework comprises four components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR) for old classes, Truncated Cross-Entropy (TCE) for new classes, and Continual Classifier Re-Training (CCRT). To tackle the issue of catastrophic forgetting, the PFGBP dynamically generates a diversity of pseudo features in an online manner, leveraging class prototypes of old classes along with batch class prototypes of new classes. Furthermore, the VPR enforces consistency between the classifier’s weights and the prototypes of old classes, leveraging class prototypes and covariance matrices to enhance robustness and generalization capabilities. The TCE mitigates the impact of domain differences of the classifier caused by pseudo features. Finally, the CCRT training strategy is designed to prevent overfitting to new classes and ensure the stability of features extracted from old classes. Extensive experiments conducted on two widely used gesture recognition datasets, namely SHREC 2017 3D and EgoGesture 3D, demonstrate that our approach outperforms existing state-of-the-art methods by 11.8% and 12.8% in terms of mean global accuracy, respectively. The code is available on https://github.com/sunao-101/PGPFR-3/.
@@ I Introduction
Hand gesture recognition is an essential research problem with broad applications in fields such as human-computer interaction, virtual reality, and assistive technology. With the development of deep learning, data-driven gesture recognition methods have made significant progress [1, 2, 3]. However, most of these methods are limited to close-set scenarios and lack the ability to handle unseen or novel gestures. In addition, they often rely on a large amount of annotated data, which is often difficult to obtain in practical applications.
@@
Data-free class-incremental learning addresses the problems of data privacy concerns and limited memory space. Typical data-free approaches use model inversion to generate pseudo samples [7, 8] or pseudo features [9, 10] of old classes, which would eliminate the need to retain a large generator and only the previously trained network. Recently, Aich et al. [8] introduce a data-free approach for class-incremental hand gesture recognition, leveraging model inversion to generate pseudo samples based on eigenvectors of the covariance matrix and the support vectors of the classifier. However, the samples obtained through model inversion for each task still need to be retained. Notably, this approach stores the inverted samples rather than the original samples, making it highly inefficient. Despite the already significant computational cost of model inversion, the requirement for vast storage to retain these inversion samples escalates as the number of continual tasks increases.
@@
Our objective is to efficiently tackle class-incremental gesture recognition using a data-free approach, as shown in  1. We focus on generating effective pseudo features for old classes, rather than pseudo samples, thereby significantly reducing computational costs. To enhance the representation of past sample distributions, we introduce class prototypes and covariance matrices for old classes. During the batch training of new class samples for each task in continual learning, pseudo features, prototypes, and new features are concurrently fed into the classifier, working together to address the issue of catastrophic forgetting.
@@
To this end, we introduce a framework called Prototype-Guided Pseudo Feature Replay (PGPFR) tailored for class-incremental gesture recognition. The PGPFR comprises four crucial components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR) for old classes, Truncated Cross-Entropy (TCE) for new classes and Continual Classifier Re-Training (CCRT). To generate a diverse set of pseudo features, the PFGBP combines class prototypes from old classes, features of new classes, and batch-specific class prototypes of new classes, which are computed within each batch of data, utilizing efficient geometric transformations. To strike a balance between overconfidence and uncertainty, temperature sharpening is used to control the steepness of probability predictions of pseudo features. The VPR is built on the assumption that the classifier ought to precisely predict the class associated with the prototype of a particular old class. A plain prototype replay loss is introduced, followed by the VPR that incorporates covariances to enhance robustness. The TCE loss aims to mitigate the influence of domain differences of the classifier [11]. Lastly, we exploit the classifier re-training strategy for class-incremental gesture recognition. It should be noted that the spatial complexity of our model is low, since the pseudo features generated by our approach are discarded immediately after being used in each batch, eliminating the need for their retention.
@@
  • We introduce an efficient Prototype-Guided Pseudo Feature Replay (PGPFR) framework for class-incremental gesture recognition. This framework is capable of continually recognizing gestures from new classes while preserving data privacy.
  • We propose a Pseudo Feature Generation with Batch Prototypes (PFGBP) module that utilizes efficient operations and batch prototypes to generate pseudo features and pseudo labels for old classes in an online fashion.
  • We propose Variational Prototype Replay (VPR) which leverages class prototypes and covariance matrices to dynamically update the classifier weights of old classes. Additionally, we introduce Continual Classifier Re-Training (CCRT) strategy to further the problem mitigate catastrophic forgetting.
@@ II-A Video-Based Hand Gesture Recognition
Hand gestures primarily consist of various combinations of fingers. Earlier works employed shape matching to measure the dissimilarity among different hand shapes. Deep learning methods learn representations directly from raw data and perform recognition in an end-to-end manner. O Köpüklü et al. [12] introduce a novel hierarchical structure for neural networks, facilitating the efficient deployment of offline-trained CNN architectures using sliding window techniques. The attention mechanism can enable the model to focus on key areas related to gesture recognition in videos or images, thereby improving recognition performance. Consequently, Dhingra et al. [13] propose an end-to-end trainable 3D attention residual neural network Res3ATN, which stacks attention blocks and can adaptively change attention features. In order to better integrate multimodal information. Elboushaki et al. [14] introduce an enhanced motion representation method using iMHI and iDMM, with multimodal fusion strategies at both feature and decision levels. Avola et al. [15] employ an enhanced stacked hourglass network for multitasking, generating 2D heatmaps and hand contours, and predicting 3D hand joint positions and meshes using the MANO layer. Recently, Li et al. [16] propose an analytical framework based on the information bottleneck theory to reduce the interference of gesture irrelevant factors by optimizing feature coding. Although these methods achieve satisfactory performance on different gesture datasets, they rely on the RGB-D video as input. This dependency causes the model susceptible to factors like background noise, illumination variations, and viewpoint changes [8], ultimately leading to a degradation in model performance.
@@
Similar to skeleton-based human action recognition [17, 18, 19], 3D hand gesture recognition refers to the understanding of hand gestures by capturing and analyzing the motion information of human joints. The 3D skeleton sequence is typically utilized as the input due to its inherent advantages in robustness and lightness. Nguyen et al. [20] introduce a novel neural network that learns the SPD matrix from skeletal data and constructs a novel layer using a variant of random gradient descent on the Stiefel manifold. De Smedt et al. [21] argue that hand skeleton data can be used to extract valid hand kinematic descriptors in gesture sequences, statistically and temporally encoded via a Fisher kernel and a multilevel temporal pyramid. Liu et al. [22] argue that gestures can be decoupled into hand pose changes and hand movements and model them separately. Analogously, Guo et al. [23] propose a new edge variation graph to divide each neighborhood of the central node into three groups of physical neighbors, temporal neighbors, and variational neighbors. Graph convolutional networks have proven to be effective in gesture recognition. Song et al. [24] propose the MS-ISTGCN, which uses a spatial and a temporal graph convolutional layer to learn the relationship between hand joints and to capture temporal features, respectively. To prevent spatial information and temporal information from influencing each other to produce fuzzy semantics, Liu et al. [25] propose TD-GCN to capture the spatiotemporal dependence between the skeleton joints by calculating the adjacency matrix of channel dependence and time dependence corresponding to different channels and frames. Transformer also plays an important role in extracting correlations between different nodes and different frames. Zhao et al. [26] propose STST that utilizes a Spatial-Temporal Synchronous Transformer to make the model focus on extracting spatial-temporal correlations simultaneously. Distinct from prior works, we focus on the challenging setting of class-incremental gesture recognition.
@@ II-C Data-Free Class-Incremental Learning
Replay-based incremental learning methods [4, 5] face issues of user data privacy and linear increase in sample storage space with the increasing number of incremental tasks. Knowledge distillation methods [27, 28, 29] use a typical student-teacher architecture to distill the knowledge of past tasks from the teacher model into the student model. To transfer past knowledge to the model under new tasks without retaining past samples or models, data-free approaches [30, 31, 32] utilize the GAN model structure to generate pseudo samples to mitigate catastrophic forgetting. The GAN model needs to be saved throughout the entire CIL process and has much larger computation and storage space compared to the classification model. Recent works use model inversion to generate past samples or features that are mixed with new class samples for model training, such as DeepInversion [7], ABD [11], and BOAT-MI [8]. Specifically, ABD [11] improves DeepInversion [7] by proposing local cross entropy for new class samples and importance-weighted feature distillation for features. BOAT-MI [8] stores the distribution information of past class samples and proposes a combination of SV-based and Proto-based methods to explore pseudo features of old classes in deep feature space. Nevertheless, the computational cost of BOAT-MI is considerable. Another disadvantage of BOAT-MI is the need to retain the pseudo samples obtained by model inversion for each task for subsequent incremental tasks, resulting in a sharp increase in the storage space of inversion samples with the increase of tasks. Our focus is data-free class-incremental gesture recognition without retaining pseudo data, and the objective is to design an efficient pseudo-feature-based approach that fully leverages the advantages of prototypes.
@@ III Methodology
For Class-Incremental Learning (CIL), the model continuously learns $N$ incremental tasks $\mathcal{T}_{0},\mathcal{T}_{1},...,\mathcal{T}_{i},...,\mathcal{T}_{N}$}. The first task $\mathcal{T}_{0}$ requires $k$ classes, and subsequently, for each incremental task , $d$ new classes $\mathcal{C}_{n}$ are added to the existing classes $\mathcal{C}_{o}$ from all previous old tasks. During each incremental task $\mathcal{T}_{i}$, a specific dataset $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$ is provided for model training. In the context of Data-Free Class-Incremental Learning (DFCIL), the model is invisible to the training set that has already been seen in the past and can only be trained on $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$. The critical issue with DFCIL is catastrophic forgetting, a phenomenon where the decision boundary of the classifier gradually becomes blurred as more incremental tasks are introduced, ultimately resulting in a substantial decline in model performance.
We propose an effective Prototype-Guided Pseudo Feature Replay (PGPFR) framework. On the one hand, we strengthen the decision boundaries of old classes by introducing Pseudo Feature Generation with Batch Prototypes (PFGBP) and Variational Prototype Replay (VPR) for old classes. On the other hand, we employ Truncated Cross-Entropy (TCE) to learn decision boundaries for new categories, alongside the Continual Classifier Re-Training (CCRT) strategy, which addresses the challenges of overfitting to instances of new classes and underfitting to those of old classes.
For convenience, we divide the model into two parts: a backbone network $f(\cdot)$ and a classifier $\varphi(\cdot)$. DG-STA [33] is employed as the backbone to effectively extract 3D gesture features, and a fully connected layer serves as the linear classifier for gesture recognition. The overall architecture of the proposed method is illustrated in Figure 2, and details are as follows.
@@ III-A Pseudo Feature Generation with Batch Prototypes
The inclusion of all new class samples in the input guarantees that pseudo features and new class samples exhibit a comparable distribution, thereby addressing the issue of data imbalance. Simple geometric translations cause the pseudo features to reside in the feature space close to the prototype of the past class, indicating the validity of pseudo features.
@@
2:  $\mathcal{F}_{\text{pseudo}}\leftarrow\{\},\mathcal{L}_{\text{pseudo}}\leftarrow\{\}$
@@
14:  $\mathcal{F}_{\text{pseudo}}\leftarrow\mathcal{F}_{\text{pseudo}}\cup\{\hat{f}(x_{p})\}$
15:  $\mathcal{L}_{\text{pseudo}}\leftarrow\mathcal{L}_{\text{pseudo}}\cup\{p\}$
@@ III-B Variational Prototype Replay for Old Classes
where $\mathcal{N}_{o}$ is a number of old classes, $w_{k}$ is the classification weight for the $k$-th class in the fully connected layer, $b_{k}$ is the corresponding bias vector, and $\mu_{k}$ is the prototype of the $k$-th class.
The original prototype lacks a measure of sample distributions of the given class. Samples closer to class prototypes in the feature space are more likely to be classified correctly, while those farther away may be misclassified, which leads to a single solution of the intra-class features. We employ the class prototypes and covariance matrices, which characterizes each class as a distribution rather than a singular point in the latent space. To enhance the robustness and generalization, we incorporate covariances and penalize the predictions of class prototypes with a regularization term in the denominator. The Variational Prototype Replay (VPR) loss is formulated as follows:

### v2

@@ Abstract
Gesture recognition is an important research area in the field of computer vision. Most existing efforts focus on close-set scenarios, thereby limiting the capacity to effectively handle unseen or novel gestures. We aim to address class-incremental gesture recognition, which entails the ability to accommodate new and previously unseen gestures over time. Specifically, we introduce a Prototype-Guided Pseudo Feature Replay framework for data-free class-incremental learning. This framework comprises four components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay for old classes, Truncated Cross-Entropy for new classes, and Continual Classifier Re-Training. To tackle the issue of catastrophic forgetting, the PFGBP dynamically generates a diversity of pseudo features in an online manner, leveraging class prototypes of old classes along with batch class prototypes of new classes. Furthermore, the Variational Prototype Replay enforces consistency between the classifier’s weights and the prototypes of old classes, leveraging class prototypes and covariance matrices to enhance robustness and generalization capabilities. The Truncated Cross-Entropy mitigates the impact of domain differences of the classifier caused by pseudo features. Finally, the Continual Classifier Re-Training training strategy is designed to prevent overfitting to new classes and ensure the stability of features extracted from old classes. Extensive experiments conducted on two widely used gesture recognition datasets, namely SHREC 2017 3D and EgoGesture 3D, demonstrate that our approach outperforms existing state-of-the-art methods by 11.8% and 12.8% in terms of mean global accuracy, respectively. The code is available on https://github.com/sunao-101/PGPFR-3/.
@@ I Introduction
Hand gesture recognition is an essential research problem with broad applications in fields such as human-computer interaction, virtual reality, and assistive technology. Recently, data-driven gesture recognition methods have made significant progress [25, 38, 22, 36]. However, most of these methods are limited to close-set scenarios and lack the ability to handle unseen or novel gestures. In addition, they often rely on a large amount of annotated data, which is often difficult to obtain in practical applications.
@@
Data-free class-incremental learning addresses the problems of data privacy concerns and limited memory space. Typical data-free approaches use model inversion to generate pseudo samples [37, 1] or pseudo features [24, 21] of old classes, which would eliminate the need to retain a large generator and only the previously trained network. Recently, Aich et al. [1] introduce a data-free approach for class-incremental hand gesture recognition, leveraging model inversion to generate pseudo samples based on eigenvectors of the covariance matrix and the support vectors of the classifier. However, the samples obtained through model inversion for each task still need to be retained. Notably, this approach stores the inverted samples rather than the original samples, making it highly inefficient. Despite the significant computational cost of model inversion, the requirement for vast storage to retain these inversion samples escalates as the number of continual tasks increases.
@@
Our objective is to efficiently tackle class-incremental gesture recognition using a data-free approach, as shown in  1. We focus on generating effective pseudo features for old classes, rather than pseudo samples, thereby significantly reducing computational costs during training. To enhance the representation of past sample distributions, we introduce class prototypes and covariance matrices for old classes. During the batch training of new class samples for each task in continual learning, pseudo features, prototypes, and new features are concurrently fed into the classifier, working together to address the issue of catastrophic forgetting.
@@
To this end, we introduce a framework called Prototype-Guided Pseudo Feature Replay (PGPFR) tailored for class-incremental gesture recognition. The PGPFR comprises four crucial components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay for old classes, Truncated Cross-Entropy for new classes and Continual Classifier Re-Training. To generate a diverse set of pseudo features, the PFGBP combines class prototypes from old classes, features of new classes, and batch-specific class prototypes of new classes, which are computed within each batch of data, utilizing efficient geometric transformations. To strike a balance between overconfidence and uncertainty, temperature sharpening is used to control the steepness of probability predictions of pseudo features. The Variational Prototype Replay is built on the assumption that the classifier ought to precisely predict the class associated with the prototype of a particular old class. A plain prototype replay loss is introduced, followed by the Variational Prototype Replay that incorporates covariances to enhance robustness. The Truncated Cross-Entropy loss aims to mitigate the influence of domain differences of the classifier [30]. Lastly, we exploit the classifier re-training strategy for class-incremental gesture recognition. It should be noted that the spatial complexity of our model is low, since the pseudo features generated by our approach are discarded immediately after being used in each batch, eliminating the need for their retention.
@@
  • To protect data privacy and improve efficiency, we introduce an efficient Prototype-Guided Pseudo Feature Replay framework for data-free class-incremental learning.
  • We propose a versatile module of Pseudo Feature Generation with Batch Prototypes that utilizes efficient operations and batch prototypes to generate pseudo features and pseudo labels for old classes in an online fashion.
  • We propose Variational Prototype Replay which leverages class prototypes and covariance matrices to dynamically update the classifier weights of old classes. Additionally, we introduce Continual Classifier Re-Training strategy to further the problem mitigate catastrophic forgetting.
@@ II-A Video-Based Hand Gesture Recognition
Hand gestures primarily consist of various combinations of fingers. Earlier works employed shape matching to measure the dissimilarity among different hand shapes. Deep learning methods learn representations directly from raw data and perform recognition in an end-to-end manner. O Köpüklü et al. [15] introduce a novel hierarchical structure for neural networks, facilitating the efficient deployment of offline-trained CNN architectures using sliding window techniques. The attention mechanism can enable the model to focus on key areas related to gesture recognition in videos or images, thereby improving recognition performance. Consequently, Dhingra et al. [9] propose an end-to-end trainable 3D attention residual neural network, which stacks attention blocks and can adaptively change attention features. In order to better integrate multimodal information, Elboushaki et al. [10] introduce an enhanced motion representation with multimodal fusion strategies at both feature and decision levels. Avola et al. [2] employ an enhanced stacked hourglass network for multitasking, generating 2D heatmaps and hand contours, and predicting 3D hand joint positions and meshes.
Recently, Li et al. [16] propose an analytical framework based on the information bottleneck theory to reduce the interference of gesture irrelevant factors by optimizing feature coding. Although these methods achieve satisfactory performance on different gesture datasets, they rely on the RGB-D video as input. This dependency causes the model susceptible to factors like background noise, illumination variations, and viewpoint changes [1], ultimately leading to a degradation in model performance.
@@
Similar to skeleton-based human action recognition [36, 35, 3, 13], 3D hand gesture recognition refers to the understanding of hand gestures by capturing and analyzing the motion information of human joints. The 3D skeleton sequence is typically utilized as the input due to its inherent advantages in robustness and lightness. Nguyen et al. [23] introduce a novel neural network that learns the positive semi-definite matrix from skeletal data and constructs a novel layer using a variant of random gradient descent on the Stiefel manifold. De Smedt et al. [7] argue that hand skeleton data can be used to extract valid hand kinematic descriptors in gesture sequences, statistically and temporally encoded via a Fisher kernel and a multilevel temporal pyramid. Liu et al. [19] argue that gestures can be decoupled into hand pose changes and hand movements and model them separately. Analogously, Guo et al. [12] propose a new edge variation graph to divide each neighborhood of the central node into three groups of physical neighbors, temporal neighbors, and variational neighbors.
Graph convolutional networks have proven to be effective in gesture recognition. Song et al. [31] use a spatial and a temporal graph convolutional layer to learn the relationship between hand joints and to capture temporal features, respectively. To prevent spatial information and temporal information from influencing each other to produce fuzzy semantics, Liu et al. [20] propose a framework to capture the spatiotemporal dependence between the skeleton joints by calculating the adjacency matrix of channel dependence and time dependence corresponding to different channels and frames. Transformer also plays an important role in extracting correlations between different nodes and different frames. Zhao et al. [40] utilize a Spatial-Temporal Synchronous Transformer to make the model focus on extracting spatial-temporal correlations simultaneously. Distinct from prior works, we focus on the challenging setting of class-incremental gesture recognition.
@@ II-C Data-Free Class-Incremental Learning
Similar to incremental clustered lifelong learning [33], generalized lifelong spectral clustering [32] and continual image generation [34, 18], data-free class-incremental learning aims to mitigate catastrophic forgetting without relying on full access to previous data. Replay-based incremental learning methods [27, 4] face issues of user data privacy and linear increase in sample storage space with the increasing number of incremental tasks. Knowledge distillation methods [17, 8, 26] use a typical student-teacher architecture to distill the knowledge of past tasks from the teacher model into the student model. To transfer past knowledge to the model under new tasks without retaining past samples or models, data-free approaches [29, 28, 6] utilize the Generative Adversarial Network (GAN) model structure to generate pseudo samples to mitigate catastrophic forgetting. The GAN model needs to be saved throughout the entire class-incremental learning process and has much larger computation and storage space compared to the classification model.
Recent works use model inversion to generate past samples or features that are mixed with new class samples for model training, such as DeepInversion [37], ABD [30], and BOAT-MI [1]. Specifically, ABD [30] improves DeepInversion [37] by proposing local cross entropy for new class samples and importance-weighted feature distillation for features. BOAT-MI [1] stores the distribution information of past class samples and proposes a combination of SV-based and Proto-based methods to explore pseudo features of old classes in deep feature space. Nevertheless, the computational cost of BOAT-MI is considerable. Another disadvantage of BOAT-MI is the need to retain the pseudo samples obtained by model inversion for each task for subsequent incremental tasks, resulting in a sharp increase in the storage space of inversion samples with the increase of tasks. Our focus is data-free class-incremental gesture recognition without retaining pseudo data, and the objective is to design an efficient pseudo-feature-based approach that fully leverages the advantages of prototypes.
@@ III Methodology
For Class-Incremental Learning, the model continuously learns $N$ incremental tasks $\mathcal{T}_{0},\mathcal{T}_{1},...,\mathcal{T}_{i},...,\mathcal{T}_{N}$}. The first task $\mathcal{T}_{0}$ requires $k$ classes, and subsequently, for each incremental task , $d$ new classes $\mathcal{C}_{n}$ are added to the existing classes $\mathcal{C}_{o}$ from all previous old tasks. During each incremental task $\mathcal{T}_{i}$, a specific dataset $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$ is provided for model training. In the context of data-free class-incremental learning, the model is invisible to the training set that has already been seen in the past and can only be trained on $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$. The critical issue with data-free class-incremental learning is catastrophic forgetting, a phenomenon where the decision boundary of the classifier gradually becomes blurred as more incremental tasks are introduced, ultimately resulting in a substantial decline in model performance.
We propose an effective Prototype-Guided Pseudo Feature Replay (PGPFR) framework. We strengthen the decision boundaries of old classes by introducing Pseudo Feature Generation with Batch Prototypes (PFGBP) and Variational Prototype Replay for old classes. In addition, we employ Truncated Cross-Entropy to learn decision boundaries for new categories, alongside the Continual Classifier Re-Training strategy, which addresses the challenges of overfitting to instances of new classes and underfitting to those of old classes.
For convenience, we divide the model into two parts: a backbone network $f(\cdot)$ and a classifier $\varphi(\cdot)$. DG-STA [5] is employed as the backbone to effectively extract 3D gesture features, and a fully connected layer serves as the linear classifier for gesture recognition. The overall architecture of the proposed method is illustrated in Figure 2.
@@ III-A Pseudo Feature Generation with Batch Prototypes
The inclusion of all new class samples in the input guarantees that pseudo features and new class samples exhibit a comparable distribution, thereby addressing the issue of data imbalance. Since the pseudo features are generated together with all new class samples, they naturally align with the same underlying feature distribution, thereby balancing the contributions of old and new classes during training. Simple geometric translations cause the pseudo features to reside in the feature space close to the prototype of the past class, indicating the validity of pseudo features.
@@
[…]
@@
[…]
@@ III-B Variational Prototype Replay for Old Classes
where $\mathcal{N}_{o}$ is a number of old classes, $w_{k}$ is the classification weight for the $k$-th class in the fully connected layer, $b_{k}$ is the bias vector, and $\mu_{k}$ is the prototype of the $k$-th class.
The original prototype lacks a measure of sample distributions of the given class. Samples closer to class prototypes in the feature space are more likely to be classified correctly, while those farther away may be misclassified, which leads to a single solution of the intra-class features. We employ the class prototypes and covariance matrices, which characterizes each class as a distribution rather than a singular point in the latent space. To enhance the robustness and generalization, we incorporate covariances and penalize the predictions of class prototypes with a regularization term in the denominator. The Variational Prototype Replay loss is formulated as follows: